How many pints of a 35% iodine solution must be added to 20 pints of a 15% iodine solution to obtain a 20% iodine solution?
Answer by stanbon(75887)

How many pints of a 35% iodine solution must be added to 20 pints of a 15% iodine solution to obtain a 20% iodine solution?
-----------------------
Equation:
iodine + iodine = iodine
0.15*20 + 0.35x = 0.20(20+x)
Multiply thru by 100 to get:
15*20 + 35x = 20*20 + 20x
15x = 5*20
x = 20/3 = 6 1/3 pints (amt. of 35% solution needed in the mixture)
=======================================================================
Cheers,
Stan H.